# GOAL: A Generalist Combinatorial Optimization Agent Learning

### v2

@@ 3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [drakulic_bq-nco_2023], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Adapters can also be introduced within layers, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used: all relational information is assumed to be captured by $\bar{x}$. The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2022, henderson_transformers_2023] but with a different design.
@@
Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as an $N{\times}D$ matrix:
@@
The score matrix is still obtained as a scalar product of a key and a query vector, but now, each of them is the sum of a node component and an edge component ((1b)). Our mixed attention allows to incorporate arbitrary relational information (edge features) into the core of the attention mechanism. It differs from positional encoding, which attempts to distill a priori the edge information (limited to sequential ordering) into the node representations.
@@
Input and output adapters are linear projections, of the task-specific instance features to the backbone’s input embedding (dim. $D$) and from the backbone’s output embedding to task-specific action space. However, when $D$ is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks. To prevent this, we simply force the input linear projection to be low rank. Concretely, node (resp. edge) features are first projected into a “small” dimension space, forcing the representations to share dimensions across tasks, then this low dimension representation is plunged into the “large” $D$-dimensional embedding space using a shared codebook (see Fig. 1 left).
@@ 3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” and “machine” nodes. For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the nodes and edges, though they are often heterogeneous by nature: in the scheduling example, edge features may carry precedence constraints when between operations, or processing times when between machines and operations, and no information at all when between machines. Instead of packing all these different types of relations into a single vector, with padding values wherever a feature is undefined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1b), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the configurations: all the attention blocks of a layer, be them self- or cross- attentions, and whatever the types they involve, share the same, layer-specific parameters. This is precisely what allows us to train a unique model to solve arbitrary single- and multi-type problems.
@@ 4 Experiments
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem (CVRP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [vidal_hybrid_2022] the Orienteering Problem (OP) with A4OP [kobeaga_efficient_2018], the Knapsack Problem (KP) with ORTools [perron_or-tools_2022], the Minimum Vertex Covering Problem (MVC) with FastWVC [cai_fastwvc_2023], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [huangfu_parallelizing_2018] and the Job Shop Scheduling Problem (JSSP) with ORTools [perron_or-tools_2022]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 100$\times$20 for UMSP (resulting in both cases in 100 operation nodes). The description of the problems and the associated data generation processes is provided in Appendix E.
Training process.
GOAL is trained by alternating batches in random order from each of the training tasks, using a single-class cross-entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross-entropy loss for KP and MVC. Similarly to [drakulic_bq-nco_2023], we create a batch by sampling suffix subproblems for each instance. We use AdamW optimizer [loshchilov_decoupled_2019], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch. We trained the model on 8 Nvidia V100 GPU servers, with batch size of 256 for 7 days, completing ca. 400 epochs. The best model was selected based on the average performance on a validation set of 128 instances per problem. The model hyperparameters are described in Appendix A.
@@
We consider the following constructive neural models as baselines: AM [kool_attention_2019], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2022], S2V-DQN [khalil_learning_2017] and COMPASS [chalumeau_combinatorial_2023]. All models except MVMoE are specialized and trained in single-task mode. They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search  [choo_simulation-guided_2022], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2022], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
@@ 4.1 Performance on the Training Tasks
In addition, in Appendix D, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances. We can note in particular the strong generalization for ATSP which seems much more challenging for the state-of-the-art baselines. This confirms the effectiveness of our mixed-attention mechanism.
@@
opt. gap time opt. gap time opt. gap time opt. gap time
@@
GOAL single-task greedy 0.21% 10s 2.31% 10s 2.79% 10s -0.35% 3s
GOAL multi-task greedy 0.50% 10s 3.06% 10s 2.86% 11s 0.28% 3s
@@
opt. gap time opt. gap time opt. gap time opt. gap time
Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 12m
@@
COMPASS - - - 4.70% 23m
GOAL single-task greedy 0.10% 3s 0.26% 3s 1.30% 7s 2.70% 15s
GOAL multi-task greedy 0.14% 2s 0.37% 9s 2.01% 7s 3.74% 15s
Table 1: Performance on 128 test instances from the training tasks. Lower is better.
@@ 4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix E). GOAL can be fine-tuned to specific problems in two modes. In both, the adapters are trained from scratch, and the backbone is also open to backpropagation (no parameter freeze). The supervised mode is identical to training, except it processes only a few (labeled) instances of the problem for a few steps. Details and results are given in Appendix B. The unsupervised mode is more challenging, but also more realistic, since we may not have access to good-quality solutions for the new problem. We propose to fine-tune the model by iterative imitation-improvement steps between an Apprentice and an Expert, as in ExIt [anthony_thinking_2017]. While Exit relies on a possibly complex tree-search for improvement, we use instead a brute-force but fast method, where the Expert samples multiple solutions (in our case 128) per instance from the Apprentice and selects the best, if it is sufficiently better, and sends it back to the Apprentice (current model) for imitation. This is similar to the self-labeling strategy proposed by corsini_self-labeling_2024 for the JSSP. One could also use more sophisticated experts such as generic local search heuristics or smarter sampling pirnay_self-improvement_2024 to accelerate the finetuning. Performance results are given in Fig 3, which shows that unsupervised fine-tuning systematically outperforms unsupervised training of the whole model from scratch at a comparable processing time budget. We also compare finetuning GOAL to finetuning the single-task models and report the results in Fig 2. We observe that the multi-task model consistently performs as well or better than the “most-related” single-task model (e.g. CVRP model for OCVRP or JSSP model for OSSP), hinting to the synergistic effect of the multi-task training.
Figure 2: Comparison of the supervised fine-tuning of the single- versus multi-task GOAL on four new problems.
@@ 4.3 Ablation Study
We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2022] and G2G [henderson_transformers_2023]. More precisely we focus on the ATSP and train three versions of GOAL with the three different variants of mixed-attention. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Fig 4 (left) presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the larger optimality gap.
@@
The proposed multi-type transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply concatenate the features of the different types of nodes (resp. edges), and use dummy “padding” values whenever a feature is undefined for a given node (resp. edge), so that the original single-type model can be used. We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 (middle) a comparison between the training performance of the single-type and multi-type architectures. Clearly, the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
@@ 5 Discussion and Conclusion
In this paper, we propose a generalist neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on seven CO problems; (ii) a multi-task model, able to solve different problems with the same backbone; and (iii) a pretrained model, which is efficiently fine-tuned to new problems.
The proposed multi-task pre-training relies on imitation learning and therefore the availability of good-quality solutions. While this is not an intrinsic limitation of the GOAL model, we believe that imitation learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [sun_smart_2023, reed_generalist_2022]. This is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research. In contrast, for a new potentially more original problem, we show that GOAL can be effectively finetuned without any supervision.
Our approach is limited to problems where solution feasibility is easy to ensure during the construction, which is a common requirement of constructive heuristics. Finally the quadratic complexity of the underlying transformer architecture makes the approach too costly for problem sizes beyond a thousand nodes. Exploring more efficient forms of attention as basis for our mixed-attention blocks would be a promising direction for better scaling.
Appendix A GOAL Hyperparameters
The GOAL backbone consists of 9 layers with an embedding size of 128, a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021]. The mixed-attention blocks have 8 heads. Low dimensions for nodes and edges are 8 and 4, respectively.

### v3

@@ 3.4 Architecture of the Model
The architecture of our model generalizes that of BQ-NCO [drakulic_bq-nco_2023], and is described in Fig. 1 left. It essentially consists of a sequence of layers whose parameters are shared by all the tasks (backbone) together with task-specific adapter modules at the input and output of the model [^1] Adapters can also be introduced within layers, but this paper considers only input-output adapters.. The input-output adapter modules for task $t$ are specific to the dimensions $F_{t},\bar{F}_{t},K_{t}$ and the meaning of the corresponding features. No positional encoding is used: all relational information is assumed to be captured by $\bar{x}$. The layers themselves, on the other hand, are only aware of the shared embedding dimensions $D,\bar{D}$. They are self-attention transformer layers [vaswani_attention_2017] where vanilla attention is replaced by multi-head mixed attention, similar in intent to [kwon_matrix_2021, henderson_transformers_2023] but with a different design.
@@
Then the key and query vectors are used to compute an $M{\times}N$ score matrix $S$, which is transformed into an attention matrix (by adding the mask and applying softmax) multiplied by the reprojected value vectors to produce the output representation $r$ of the $N$ input query nodes as a matrix with $N$ rows:
@@
The score matrix is still obtained as a scalar product of a key and a query vector, but now, each of them is the sum of a node component and an edge component ((1b)). Our mixed attention allows to incorporate arbitrary relational information (edge features) into the core of the attention mechanism. It differs from positional encoding, which attempts to instill a priori the edge information (limited to sequential ordering) into the node representations.
@@
Input and output adapters are linear projections, from the task-specific instance features to the backbone’s input embeddings and from the backbone’s output embeddings to the task-specific action space. However, when the embedding dimension is large and the number of training tasks is small, there is a risk that the embeddings produced by each task end up occupying complementary subspaces of the embedding space throughout the layers, which would defeat the purpose of effectively sharing parameters across tasks. To prevent this, we simply force the input linear projection to be low rank. Thus, node (resp. edge) features are first mapped into a low dimension $\ell{\ll}D$ (resp. ¯$\bar{\ell}{\ll}\bar{D}$), through a task-specific linear projection (the input adapter proper), forcing the representations to share dimensions across tasks. Then this “small” representation is plunged into a “large” $D$- (resp. $\bar{D}$)-dimensional embedding, through a common linear projection, called a codebook, shared by all the tasks (see Fig. 1 left).
@@ 3.5 Multi-Type Transformer Architecture
Many tasks (CO problems) involve different types of nodes. For example, scheduling problems often involve “operation” and “machine” nodes. For GOAL, one solution is to simply treat them all indiscriminately in the backbone. The adapters, being problem specific, are of course free to treat them differently according to their types. This is sometimes unsatisfactory though, since it forces the adapter to provide uniform embeddings for all the nodes and edges, though they are often heterogeneous by nature: in the scheduling example, edge features may carry precedence constraints when between operations, or processing times when between machines and operations, and no information at all when between machines. Instead of packing all these different types of relations into a single vector, with padding values wherever a feature is undefined for a given edge, we propose a multi-type architecture where the adapters output separate embeddings per type (for nodes) or compatible pairs of types (for edges). Each layer of the model, instead of being a monolithic self-attention block operating on all the nodes, becomes a task-specific combination of both cross- and self-attention blocks, as described by (1b), operating on different type-compatible subsets of nodes and edges. This is illustrated in Fig. 1 right in the case of two types. For $n$ types, the architecture would be similar but involves $n^{2}$ mixed-attention blocks, where each type attends to itself and every other type, or a selection of relevant types defined by the task. Therefore, in our scheduling example, the model will contain separate self-attention blocks, for operation nodes and machine nodes respectively, and cross-attention blocks, for operation-machine edges and vice-versa. Observe that while the architecture configuration is now task-specific, the model parameters remain the same for all the configurations: all the attention blocks of a layer, be them self- or cross- attentions, and whatever the types they involve, share the same, layer-specific parameters. This is precisely what allows us to train a unique model to solve arbitrary single- and multi-type problems.
3.6 Training Procedure
We train GOAL on a set of eight various CO problems (described in Sec 4). For each training problem, we define the corresponding BQ-MDP, the linear input and output adapters to match the dimensions of the input features and the construction actions. We train the model by imitation of high-quality solutions, provided by specialized solvers. At each step, we sample a problem and create a batch of instances by sampling suffix subproblems from the training instances, similarly to [drakulic_bq-nco_2023]. The forward and backward processes involve the input and output adapters, as well the backbone parameters, where the mixed-attention and feed-forward blocks are duplicated within each layer if the sampled task is multi-typed.
@@ 4 Experiments
We train GOAL on eight classic and varied CO problems. For each problem, we generate a training dataset of 1 million random instances and solve them using traditional solvers. Specifically, we consider the following problems and oracle solvers: the Asymmetric Traveling Salesman Problem (ATSP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem (CVRP) with LKH [helsgaun_extension_2017], the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) with HGS [vidal_hybrid_2022] the Orienteering Problem (OP) with A4OP [kobeaga_efficient_2018], the Knapsack Problem (KP) with ORTools [perron_or-tools_2022], the Minimum Vertex Covering Problem (MVC) with FastWVC [cai_fastwvc_2023], the Unrelated Machine Scheduling Problem (UMSP) with HiGHS [huangfu_parallelizing_2018] and the Job Shop Scheduling Problem (JSSP) with ORTools [perron_or-tools_2022]. The training instances are of size 100 nodes, or 10$\times$10 for JSSP and 100$\times$20 for UMSP (resulting in both cases in 100 operation nodes). The description of the problems and the associated data generation processes is provided in Appendix F.
Training details.
GOAL is trained by alternating batches in random order from each of the training tasks, using a single-class cross-entropy loss for ATSP, CVRP, CVRPTW, OP, JSSP and UMSP and a multi-class cross-entropy loss for KP and MVC. We use AdamW optimizer [loshchilov_decoupled_2018], with an initial learning rate of 0.0005 and a decay rate of 0.97 every 10th epoch. We trained the model on 8 Nvidia V100 GPU servers, with batch size of 256 for 7 days, completing ca. 400 epochs. The best model was selected based on the average performance on a validation set of 128 instances per problem. The model hyperparameters are described in Appendix A.
@@
We consider the following constructive neural models as baselines: AM [kool_attention_2018], MDAM [xin_multi-decoder_2021], POMO [kwon_pomo_2020], Sym-NCO [kim_sym-nco_2022], BQ-NCO [drakulic_bq-nco_2023], MVMoE [zhou_mvmoe_2024], MatNet [kwon_matrix_2021], S2V-DQN [khalil_learning_2017], RouteFinder [berto_routefinder_2024], COMPASS [chalumeau_combinatorial_2023] and Gumbeldore [pirnay_self-improvement_2024]. All models except MVMoE and RouteFinder are specialized and trained in single-task mode. They are often combined at test time with various enhancers, such as beam search [drakulic_bq-nco_2023], simulation-guided beam search [choo_simulation-guided_2022], MCTS [xing_graph_2020], instance augmentation [kwon_pomo_2020], or active search [hottung_efficient_2021], but for fair comparison, all the results we report are obtained by a pure greedy application of the trained policy at test time.
@@ 4.1 Performance on the Training Tasks
In addition, in Appendix E, we provide experimental results demonstrating GOAL’s excellent zero-shot generalization on instances up to 10 times larger than the training instances. We can note in particular the strong generalization for ATSP which seems much more challenging for the state-of-the-art baselines. This confirms the effectiveness of our mixed-attention mechanism.
@@
gap time gap time gap time gap time
@@
GOAL single-task greedy 0.30% 10s 2.34% 10s 2.61% 10s -0.04% 3s
GOAL multi-task greedy 0.91% 10s 3.16% 10s 3.82% 10s 0.43% 3s
@@
gap time gap time gap time gap time
Oracle solver 0.00% 1s 0.00% 2m 0.00% 2m 0.00% 47s
@@
COMPASS - - - 4.70% 3h
Gumbeldore greedy - - - 3.17% 9s
GOAL single-task greedy 0.10% 3s 0.23% 3s 2.82% 7s 2.73% 15s
GOAL multi-task greedy 0.12% 3s 0.37% 3s 3.84% 7s 4.13% 15s
Table 1: Performance on 128 test instances from the training tasks. Each gap is with respect to the oracle and often corresponds to the optimality gap. Lower is better.
@@ 4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider eight diverse new COPs: the Traveling Repairman Problem (TRP), the Prize Collecting TSP (PCTSP), the Open CVRP (OCVRP), the Split Delivery CVRP (SDCVRP), the Sequential Ordering Problem (SOP), the Maximum Covering Location Problem (MCLP), the Maximal Independent Set (MIS) problem and the Open Shop Scheduling Problem (OSSP) (detailed descriptions in Appendix F). GOAL can be fine-tuned to specific problems in two modes. In both, the adapters are trained from scratch, and the backbone is also open to backpropagation (no parameter freeze). The supervised mode is identical to training, except it processes only a few (labeled) instances of the problem for a few steps. Details and results are given in Appendix B. The unsupervised mode is more challenging, but also more realistic, since we may not have access to good-quality solutions for the new problem. We propose to fine-tune the model by iterative imitation-improvement steps between an Apprentice and an Expert, as in ExIt [anthony_thinking_2017]. While Exit relies on a possibly complex tree-search for improvement, we use instead a brute-force but fast method, where the Expert samples multiple solutions (in our case 128) per instance from the Apprentice and selects the best, if it is sufficiently better, and sends it back to the Apprentice (current model) for imitation. This is similar to the self-labeling strategy proposed by corsini_self-labeling_2024 for the JSSP. One could also use more sophisticated experts such as generic local search heuristics or smarter sampling [pirnay_self-improvement_2024] to accelerate the finetuning. Performance results are given in Fig 2, which shows that unsupervised fine-tuning systematically outperforms unsupervised training of the whole model from scratch at a comparable processing time budget. We also compare finetuning GOAL to finetuning the single-task models and report the results in Fig 4. We observe that the multi-task model consistently performs as well or better than the “most-related” single-task model (e.g. CVRP model for OCVRP or JSSP model for OSSP), hinting to the synergistic effect of the multi-task training.
@@ 4.3 Ablation Study
We compare our mixed-attention blocks with alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [kwon_matrix_2021] and G2G [henderson_transformers_2023]. More precisely we focus on the ATSP and train three versions of GOAL with the three different variants of mixed-attention. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Fig 3 (left) presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the larger optimality gap.
@@
Our multi-type architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where job and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node and edge heterogeneity: we simply gather in a single tensor the node (edge) representations of the different types, produced by the input adapter, and use the original single-type model. Note that because of the linearity of the input adapters, these representations can be obtained either by defining a specific linear transform per type, or equivalently by defining a single linear transform on the direct sum of the feature spaces of the different types. We implemented the former for the Job Shop Scheduling Problem and present in Fig 3 (middle) a comparison between the training performance of the single-type and multi-type architectures. Clearly, the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%. Finally, note that the multi-type model offers more flexibility for fine-tuning: while, in order to be task agnostic at training, parameter sharing between the type specific attention blocks of a layer must be enforced, that constraint can be relaxed at fine-tuning to better fit the task at hand.
@@
Figure 4: Comparison of the (supervised) fine-tuning of the single- versus multi-task GOAL on four new problems.
@@ 5 Discussion and Conclusion
In this paper, we propose a generalist neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on seven CO problems; (ii) a multi-task model, able to solve different problems with the same backbone; and (iii) a pretrained model, which is efficiently fine-tuned to new problems. The proposed multi-task pre-training relies on imitation learning and therefore the availability of good-quality solutions. While this is not an intrinsic limitation of the GOAL model, we believe that imitation learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [sun_smart_2023, reed_generalist_2022]. It is particularly advantageous for CO as we can leverage high-quality solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research. In contrast, for a new potentially more original problem, we show that GOAL can be effectively finetuned without any supervision. Our approach is limited to problems where solution feasibility is easy to ensure during the construction, which is a common requirement of constructive heuristics. Finally the quadratic complexity of the underlying transformer architecture makes the approach too costly for problem sizes beyond a thousand nodes. Exploring more efficient forms of attention as basis for our mixed-attention blocks would be a promising direction for better scaling.
Appendix A GOAL Hyperparameters and size
The GOAL backbone consists of $L{=}9$ layers with an embedding size of $D{=}\bar{D}{=}128$ (the same for both nodes and edges), a feed-forward dimension of 512, and ReZero normalization [bachlechner_rezero_2021]. The mixed-attention blocks have 8 heads. The task specific input adapters produce representations of low dimension $\ell{=}8$ for nodes and ¯$\bar{\ell}{=}4$ for edges, which are mapped to embeddings by a shared codebook of dimension ×$\ell{\times}D$ for nodes and ¯×$\bar{\ell}{\times}\bar{D}$ for edges, respectively. GOAL backbone has 2.1M parameters and adapters are very light, with only a few thousand parameters per task. For the illustration, RouteFinder [berto_routefinder_2024] with POMO has 1.3M, MVMoE [zhou_mvmoe_2024] has 3.7M (the same model is used in RouteFinder with MVMoE), and RouteFinder with the Transformer Encoder 1.7M.